OpenStudy (anonymous):

I. Let f(x) = 2x – 1, g(x) = 3x, and h(x) = x2 + 1. Compute the following: f(h)(7))

OpenStudy (shaik0124):

ho(gof)(x) ho(g(f(X)) ho(g(2x-1) ho(3(2x-1) g(x)=3x and x=2x-1 so g(x)=3(2x-1) h(6x-3) h(6(x^2+1)-3) h(6x^2+6-3) therefore h(7)=6*7^2+3 =6*7*7+3 solve u will get answer